# Gaussian Control with Hierarchical Semantic Graphs in 3D Human Recovery

### v1

@@ 4.1 Datasets and Metrics
ZJU-MoCap Dataset. ZJU-Mocap [35] is a prominent benchmark in human modeling from videos. Similar to previous work [8, 43], we select 6 human subjects (377, 386, 387, 392, 393, 394) from the dataset to conduct experiments. Following [8], we used one camera for training and the remaining cameras for evaluation. For each subject, we sample 1 frame every 5 frames and collect 100 frames for training.
MonoCap Dataset. The MonoCap Dataset includes four multi-view videos from DeepCap [11] and DynaCap [12] datasets, collected by [36]. This dataset provides essential details like camera parameters, SMPL parameters, and human masks. Same as [36], we used one camera view for training and selected ten uniformly distributed cameras for testing. For each subject, we sample 1 frame every 5 frames and collect 100 frames for training.
Evaluation Metrics. We used PSNR, SSIM, and LPIPS as quality evaluation metrics. The average results of the 6 sub-datasets of the ZJU-MoCap dataset were used as the experimental results for each method on this dataset. Similarly, the average results of the 4 sub-datasets of the MonoCap dataset were used as the experimental results for each method on this dataset.
@@ 4.2 Quantitative Results
To verify the effectiveness of our method in solving the geometric distortion problem in reconstructing the human body, we compared our method with NeuralBody [35], HumanNeRF [43] AnimateNeRF [34], InstantNVR [8] InstantAvatar [18] GauHuman [14] on the ZJU dataset and the Monocap dataset, as shown in Table 1. Compared with previous methods, Our model achieved the best results on both PSNR and LPIPS on both datasets, outperforming previous methods.
@@
Table 2 shows the results of the ablation experiment on parameter k mentioned in Equation 7, showing that our model performs the best when $k$=3. This parameter is the number of different clusters of the human topological graph.
Table 2: Quantitative Results of parameter $k$
$k$ PSNR↑ SSIM↑ LPIPS↓
2 31.54 0.9681 0.01756
3 32.16 0.9758 0.01738
4 31.86 0.9673 0.01751
5 31.37 0.9632 0.01764
@@ 4.3 Qualitative Results
To address the geometric distortion stemming from semantic ambiguities at the junction of body parts, we conducted a comparative analysis of the reconstruction outcomes between the baseline method and our proposed Semantic Kinematic Topology approach across three distinct scenarios illustrated in Figure 4. The experimental findings reveal that our model outperforms the baseline methods by producing superior results not only at the intersections of human joints and clothing but also at the junctures within the human body, demonstrating clear edge boundaries. Additionally, our model exhibits improved capability in managing geometric distortions and accurately restoring the shapes of body parts in the presence of occlusions. Among them, InstantNVR [8] proposes to decompose the human body into multiple structured geometries, but only based on hierarchical shapes and textures for partitioning, without considering rich human semantic priors and lacking understanding of human motion characteristics. Although Gauhuman [14] introduced the human prior of SMPL, there is a lack of learning about topological relationships between different body parts, resulting in blurred joint junctions. AnimateNeRF [34] introduces the human joint skeleton as a regularized learning mechanism for the linkage effects between bones, but lacks precise modeling of the linkage effects between specific body parts and other body parts, resulting in the problem of human arm deformities in the rendering results. Based on learning Gaussian human body semantics, our method learns the topological relationships of different body parts, thereby deeply learning the kinematic characteristics of different body parts. This makes our method more advantageous in dealing with semantic confusion caused by complex motion postures and can generate more accurate and natural human body reconstruction results.
@@ 4.4 Ablation Study
To verify the effectiveness of our SAG (Semantic Alignment Gaussian) module, IJG (Inter Joint Graph) module, and SD(Surface Disentanglement) module, we conduct ablation experiments on the ZJU dataset and Monocap dataset. The quantitative analysis obtained is shown in Table 3.
Semantic Alignment Gaussian We conduct ablation experiments on the 377 sequence of the ZJU-Mocap dataset. The qualitative results are shown in the Figure 3. It can be seen that before adding SAG, the Gaussian points cannot be correctly matched to the corresponding geometric spatial positions of the human body parts, resulting in local geometric structure distortion in the rendering results and blurring between parts, as indicated by the boxes in Figure 3. After adding the SAG module, our model injects human body structure information into the 3D Gaussian optimization process, explicitly indicating specific body part semantics for Gaussian points in the 3D space, so that each body part can independently guide the corresponding Gaussian points for optimization, resulting in clearer rendering results for body parts.
@@
Inter-Joint Graph: Comparing the model with both SAG and IJG against the model with only SAG, as depicted in Figure 3, our IJG model utilizes a semantic topology graph to leverage the implicit topological relationships of human body parts on surface Gaussian points, resulting in a more accurate reconstruction of the human body. Notably, the representation of the geometric structure at the knees, as indicated by the blue box in Figure 3, aligns more closely with the ground truth when employing the IJG model.
Surface Disentanglement: Ablation experiments were conducted on the Surface Disentanglement (SD) module to showcase its impact on high-frequency features. The results are presented in Table 3. A comparison with the model lacking an SD module demonstrates that the complete model can effectively capture high-frequency information on the human body surface. By performing dense operations on Gaussian points in these regions to guide density and attribute optimization, the model enhances the local rendering granularity of the human body surface, presenting high-frequency details more accurately.
To further demonstrate the optimization of the SD module in reconstructing high-frequency features of the human body, we use the Canny algorithm and Fourier transform to extract the high-frequency regions of the experimental results, which are shown in Figure 5. The model lacking the Surface Disentanglement (SD) module exhibits inadequate responsiveness to high-frequency parts, resulting in incomplete reconstruction of the geometric structure for parts with significant high-frequency information, such as the face and clothing wrinkles. On the other hand, our complete model demonstrates a higher level of restoration for high-frequency features. The SD module performs effectively in restoring high-frequency surface positions, generating more realistic and stable results. This validates the efficacy of the SD module in controlling the density of Gaussian points on the high-frequency surface and optimizing their attributes.
@@ 4.5 T-Test
We conducted ten experiments on the 377 sequences of ZJU-Mocap dataset, with each experiment using a different camera for training and the remainder for validation. Identical experiments were conducted on both GauHuman and our method because GauHuman ranks second in both PSNR and LPIPS indicators (Table 1). Then through a T-Test, we validated that our method outperforms on both PSNR and LPIPS metrics (Table 4). We set the significance level at 0.05, where a p-value below this threshold indicates a difference in results between the two methods. The critical value of the T-Test, with 18 degrees of freedom, was determined to be 2.101.
We find that our method performs better in terms of PSNR, with a t-value of 2.444 (greater than the critical value of 2.101), and a p-value<0.05, indicating the correctness of the result. Similarly, on LPIPS, the t-value exceeds 2.101, and the p-value<0.05, confirming our method’s excellent performance.
@@ 5 Conclusion
We introduce a method for Gaussian human reconstruction in semantic and topological spaces to mitigate the semantic ambiguities arising from body part occlusion during human motion. Our approach explicitly integrates kinematic correlations to reconstruct more intricate human poses and deformations. Through extensive experimentation, we have validated the efficacy of our proposed method, particularly in capturing surface details and the articulation of body parts. We believe that this innovative approach can provide insights into resolving longstanding reconstruction challenges. For instance, by integrating semantic and topological information across temporal sequences to capture the evolving semantic and topological trends in human motion, enabling the utilization of 3DGS in 4D human dynamic modeling. Furthermore, we aim to explore the interplay of topological and semantic relationships between humans and objects to enhance the precision of human reconstruction in complex multi-object interaction scenarios.

### v3

@@ 4.1 Datasets and Metrics
ZJU-MoCap Dataset. The ZJU-Mocap dataset [35] is a prominent benchmark in human modeling from videos. Similar to previous work [5, 43], we select 6 human subjects (377, 386, 387, 392, 393, 394) from the dataset to conduct experiments. Following [5], we use one camera for training and the remaining cameras for evaluation. For each subject, we sample 1 frame every 5 frames and collect 100 frames for training.
MonoCap Dataset. The MonoCap Dataset includes four multi-view videos from DeepCap [8] and DynaCap [9] datasets, collected by [36]. This dataset provides essential details like camera parameters, SMPL parameters, and human masks. As in [36], we use one camera view for training and selected ten uniformly distributed cameras for testing. For each subject, we sample 1 frame every 5 frames and collect 100 frames for training.
Evaluation Metrics. We use PSNR, SSIM, and LPIPS as quality evaluation metrics. The average results of the 6 sub-datasets of the ZJU-MoCap dataset are used as the experimental results for each method on this dataset. Similarly, the average results of the 4 sub-datasets of the MonoCap dataset are used as the experimental results for each method on this dataset.
@@ 4.2 Quantitative Results
To verify the effectiveness of our method in solving the geometric distortion problem in reconstructing the human body, we compare our method with NeuralBody [35], HumanNeRF [43] AnimateNeRF [34], InstantNVR [5] InstantAvatar [15] GauHuman [11] on the ZJU dataset and the Monocap dataset, as shown in Table 1. Compared with previous methods, our model achieve the best results on both PSNR and LPIPS on both datasets, outperforming previous methods.
@@ 4.3 Qualitative Results
To address the geometric distortion stemming from semantic ambiguities at the junction of body parts, we conducted a comparative analysis of the reconstruction outcomes between the baseline method and our proposed Semantic Kinematic Topology approach across three distinct scenarios illustrated in Figure 4. The experimental findings reveal that our model outperforms the baseline methods by producing superior results not only at the intersections of human joints and clothing but also at the junctures within the human body, demonstrating clear edge boundaries. Additionally, our model exhibits improved capability in managing geometric distortions and accurately restoring the shapes of body parts in the presence of occlusions. Among the compared methods, InstantNVR [5] proposes to decompose the human body into multiple structured geometries, but it relies solely on hierarchical shapes and textures for partitioning, without considering rich human semantic priors and lacking an understanding of human motion characteristics. Although Gauhuman [11] introduced the human prior of SMPL, it fails to learn the topological relationships between different body parts, resulting in blurred joint junctions. AnimateNeRF [34] introduces the human joint skeleton as a regularized learning mechanism for the linkage effects between bones, but lacks precise modeling of the linkage effects between specific body parts and other body parts, leading to human arm deformities in the rendering results. By learning Gaussian human body semantics, our method captures the topological relationships of different body parts, thereby deeply understanding the kinematic characteristics of these parts. This makes our method more advantageous in dealing with semantic confusion caused by complex motion postures and capable of generating more accurate and natural human body reconstruction results.
@@ 4.4 Ablation Study
To verify the effectiveness of our SAG (Semantic Alignment Gaussian) module, IJG (Inter Joint Graph) module, and SD(Surface Disentanglement) module, we perform ablation experiments on the ZJU dataset and Monocap dataset. The quantitative results are shown in Table 2.
Semantic Alignment Gaussian. We conduct ablation experiments on the 377 sequence of the ZJU-Mocap dataset. The qualitative results are shown in the Figure 3. As shown in Figure 3, before adding SAG, the Gaussian points fail to be correctly matched to the corresponding geometric spatial positions of the human body parts, resulting in local geometric structure distortion in the rendering results and blurring between parts, as indicated by the boxes in Figure 3. After adding the SAG module, our model incorporates human body structure information into the 3D Gaussian optimization process, explicitly indicating specific body part semantics for Gaussian points in the 3D space, so that each body part can independently guide the corresponding Gaussian points for optimization, resulting in clearer rendering results for body parts.
@@
To validate the optimal number of categories in the k-nearest neighbors clustering algorithm, we conducted ablation experiments. Table 4 shows the results of the ablation experiment on parameter $k$ mentioned in Equation 3. This parameter is the number of different clusters of the human topological graph. Table 4shows that our model performs the best when $k$=3.
Table 4: Quantitative Results of parameter $k$
$k$ PSNR↑ SSIM↑ LPIPS↓
2 31.54 0.9681 0.01756
3 32.16 0.9758 0.01738
4 31.86 0.9673 0.01751
5 31.37 0.9632 0.01764
@@
Inter-Joint Graph. Comparing the model with both SAG and IJG against the model with only SAG, as depicted in Figure 3, our IJG model employs a semantic topology graph to leverage the implicit topological relationships of human body parts on surface Gaussian points, resulting in a more accurate reconstruction of the human body. Notably, the representation of the geometric structure at the knees, as indicated by the blue box in Figure 3, aligns more closely with the ground truth when employing the IJG model.
Surface Disentanglement: Ablation experiments were conducted on the Surface Disentanglement (SD) module to showcase its impact on high-frequency features. The results are presented in Table 2. A comparison with the model without the SD module demonstrates that the complete model can effectively capture high-frequency information on the human body surface. By performing dense operations on Gaussian points in these regions to guide density and attribute optimization, the model enhances the local rendering granularity of the human body surface, presenting high-frequency details more accurately.
To further demonstrate the optimization of the SD module in reconstructing high-frequency features of the human body, we apply the Canny algorithm and Fourier transform to extract the high-frequency regions of the experimental results, which are shown in Figure 5. The model lacking the Surface Disentanglement (SD) module exhibits inadequate responsiveness to high-frequency parts, resulting in incomplete reconstruction of the geometric structure for parts with significant high-frequency information, such as the face and clothing wrinkles. On the other hand, our complete model demonstrates a higher level of restoration for high-frequency features. The SD module performs effectively in restoring high-frequency surface positions, generating more realistic and stable results. This validates the efficacy of the SD module in controlling the density of Gaussian points on the high-frequency surface and optimizing their attributes.
@@ 4.5 T-Test
We perform ten experiments on the 377 sequences of the ZJU-Mocap dataset, with each experiment using a different camera for training and the remainder for validation. Identical experiments are performed on both GauHuman and our method because GauHuman ranks second in both PSNR and LPIPS indicators (Table 1). Through a T-Test, we validate that our method outperforms GauHuman on both PSNR and LPIPS metrics (Table 3). We set the significance level at 0.05, where a p-value below this threshold indicates a difference in results between the two methods. The critical value of the T-Test, with 18 degrees of freedom, is determined to be 2.101.
We find that our method performs better in terms of PSNR, with a t-value of 2.444 (greater than the critical value of 2.101), and a p-value<0.05, indicating the correctness of the result. Similarly, on LPIPS, the t-value also exceeds 2.101, and the p-value<0.05, confirming our method’s excellent performance.
@@ 5 Conclusion
We introduce a Human Gaussian Control with Hierarchical Semantic Graphs (HUGS) method for human reconstruction. This method utilizes the Inter-Semantic Kinematic Topology Aware module to learn the relationships in motion at the connections of different body parts and the Intro-Surface Disentanglement module to learn the relationships in appearance within body parts. It addresses blurred details in the reconstruction of the human body, at the junctions and surface features. Our approach explicitly integrates relationships in motion to reconstruct more intricate human poses and deformations. Through extensive experimentation, we have validated the efficacy of our proposed method, particularly in capturing surface details and articulating body parts. We believe that this innovative approach can provide insights into resolving longstanding reconstruction challenges. Specifically, by integrating semantic and topological information across temporal sequences, it aims to capture the evolving semantic and topological trends in human motion, thereby enabling the utilization of 3DGS in 4D human dynamic modeling. Furthermore, we aim to explore the interplay of topological and semantic relationships between humans and objects to enhance the precision of human reconstruction in complex multi-object interaction scenarios.